# Softmax Attention with Constant Cost per Token

### v1

@@ Abstract
We propose a simple modification to the conventional attention mechanism applied by Transformers: Instead of quantifying pairwise query-key similarity with scaled dot-products, we quantify it with the logarithms of scaled dot-products of exponentials. Attention becomes expressible as a composition of log-sums of exponentials that is linearizable, with a latent space of constant size, enabling sequential application with constant time and space complexity per token. We implement our modification, verify that it works in practice, and conclude that it is a promising alternative to conventional attention. [^1] Source code and instructions for replicating our results are online at https://github.com/glassroom/heinsen_attention.
@@ 1 Summary
Numerous alternatives to conventional attention have been proposed to reduce its quadratic cost, including linearized, low-rank, and sparse approximations, mechanisms that slide context windows, and convolutions in the frequency domain. We cannot summarize all proposals fairly in the space of a paragraph. Notable examples include the methods proposed by Child et al. (2019), Wang et al. (2020), Kitaev et al. (2020), Katharopoulos et al. (2020), Zhai et al. (2021), Roy et al. (2020), Lee-Thorp et al. (2021), and Poli et al. (2023). More recently, generalized state space models that build on previous research Martin and Cundy (2017) Gu et al. (2021) have shown promise by incorporating data-driven mechanisms to control the evolution of a fixed-size latent state Peng et al. (2023) Gu and Dao (2023) Katsch (2023), but their performance is inferior on certain tasks (e.g., recalling arbitrary parts of the input context), motivating the hypothesis that methods with a fixed-size latent space cannot outperform conventional attention’s quadratic-cost application of a Softmax function over the rows of a matrix with pairwise similarity scores Jelassi et al. (2024).
@@ 1.1 Modifying Attention
We find that a simple modification to conventional attention enables its linearization as a composition of log-sums of exponentials, with a fixed-size latent space, for sequential application with constant cost per token. We implement our modification, verify that it works, and conclude that it is a promising alternative to conventional attention.
@@
where queries $Q$, keys $K$ and values $V$ have $n_{Q}\times d_{K}$, $n_{K}\times d_{K}$, and $n_{K}\times d_{V}$ elements, respectively, and $c$ is a scalar constant, all in $\mathbb{R}$. We compute all exponentials elementwise. Each exponentiated query and each exponentiated key is in $\mathbb{R}_{+}^{d_{K}}$, so their scaled dot-product is in $\mathbb{R}_{+}$, and its logarithm is a conventional logit in $\mathbb{R}$.
1.2 Linearizing Attention
@@ 2 Proof
For clarity’s sake, we walk step-by-step through a sequence of algebraic manipulations. We start by expanding the Softmax function in (1) and simplifying the resulting expression. We obtain:
@@
where $\sum_{[n_{K}]}$ normalizes each row to a probability distribution. The scaling constant $\exp(c)$ disappears because it becomes a common divisor of numerator and denominator expressions.

### v2

@@ Abstract
We propose a simple modification to the conventional attention mechanism applied by Transformers: Instead of quantifying pairwise query-key similarity with scaled dot-products, we quantify it with the logarithms of scaled dot-products of exponentials. Our modification linearizes attention with exponential kernel feature maps, whose corresponding feature function is infinite dimensional. We show that our modification is expressible as a composition of log-sums of exponentials, with a latent space of constant size, enabling application with constant time and space complexity per token. We implement our modification, verify that it works in practice, and conclude that it is a promising alternative to conventional attention. [^1] Source code and instructions for replicating our results are online at https://github.com/glassroom/heinsen_attention.
@@ 1 Summary
Numerous alternatives to conventional attention have been proposed to reduce its quadratic cost, including linearized, low-rank, and sparse approximations, mechanisms that slide context windows, and convolutions in the frequency domain. We cannot summarize all proposals fairly in the space of a paragraph. Notable examples include the methods proposed by Child et al. (2019), Wang et al. (2020), Kitaev et al. (2020), Katharopoulos et al. (2020), Zhai et al. (2021), Roy et al. (2020), Lee-Thorp et al. (2021), and Poli et al. (2023). More recently, generalized state space models that build on previous research Martin and Cundy (2017) Gu et al. (2021) have shown promise by incorporating data-driven mechanisms to control the evolution of a fixed-size latent state Peng et al. (2023) Gu and Dao (2023) Katsch (2023), but their performance is inferior on certain tasks (e.g., recalling arbitrary parts of the input context), motivating the hypothesis that methods with a fixed-size latent space cannot outperform conventional attention Jelassi et al. (2024).
@@ 1.1 Modifying Attention
We find that a simple modification to conventional attention linearizes it Katharopoulos et al. (2020) with exponential kernel feature maps, and we show that this modification renders attention expressible as a composition of log-sums of exponentials, with a fixed-size latent space, for sequential application with constant cost per token. We implement our modification, verify that it works, and conclude that it is a promising alternative.
@@
where queries $Q$, keys $K$ and values $V$ have $n_{Q}\times d_{K}$, $n_{K}\times d_{K}$, and $n_{K}\times d_{V}$ elements, respectively, and $c$ is a scalar constant, all in $\mathbb{R}$. We compute all exponentials elementwise.
1.2 As Log-Sums of Exponentials
@@ 2 Proof
For clarity’s sake, we walk step-by-step through a sequence of algebraic manipulations. We start by expanding the Softmax function in (1) and simplifying the resulting expression. We obtain a form of linear attention Katharopoulos et al. (2020) with exponential kernel feature maps:
@@
where $\sum_{[n_{K}]}$ normalizes each row to a probability distribution. The scaling constant $\exp(c)$ disappears because it becomes a common divisor of numerator and denominator expressions. Note that the feature function corresponding to the exponential kernel is infinite dimensional.